We know that the yen and the swiss franc have a 120yen/ sf 1 exchange rate, meaning one swiss franc buys 120 yen in the spot ER market. If the swiss franc has an interest rate of .06 and the yen rate is -.02, what is the forward exchange rate for IPT (interest parity theory) to be attained? Show everything in yen terms, i., e., how much yen one Swiss franc buys (yen is in the numerator.) If there is no equilibrium initially, will there be equilibrium eventually? If so, what will transpire? Be extremely thorough. Your answer should include CIA.
Answer #1

Details provided : 1 Swiss Franc = 120 yen Swiss franc interest rate = 6% and yen rate = -2%

As per interest parity theory, forward rate is given by

F = So * (1+ic)/(1+ip)

F = 120 * ((1+0.06)/(1-0.02)) = 129.8

Hence one swiss franc will buy 129.8 yen in future

Even if there is no equilibrium initially it will prevail eventually as arbitrators will take advantage of imbalance to the point where there is no more arbitrage